OpenStudy (anonymous):

How do you solve 2s with 3 squared - 4r with 2 squared if r =1 and s = 5?

OpenStudy (sasogeek):

your question isn't clear enough. please write the equation you're trying to solve mathematically.

OpenStudy (anonymous):



OpenStudy (anonymous):

r =1 and s = 5?

OpenStudy (anonymous):

2(5)^3-4(1)2 = 250 - 4 = 246

OpenStudy (anonymous):

Thank you so much :)